89 octane gas is made by mixing 87 octane gas and 92 octane gas. How many gallons of the 87 octane gas and 92 octane gas should be used to make 500 gallons of the 89 octane gas?

Answer by stanbon(75887)

89 octane gas is made by mixing 87 octane gas and 92 octane gas. How many gallons of the 87 octane gas and 92 octane gas should be used to make 500 gallons of the 89 octane gas?
--------------------------------
Equation:
octane + octane = octane
87x + 92(500-x) = 89*500
87x = 92*500 - 92x = 89*500
-5x = -3*500
x = 300 (# of gallons of 87 octane gas needed in the mixture)
----------------------
500-x = 200 (# of gallons of 92 octain gas needed in the mixture)
====================================================================
Cheers,
Stan H.